### v5

@@ 4.1 The SDP relaxation
Here, we provide a brief summary of the SDP relaxation and its properties. See [1] for a more detailed description.
@@
subject to $\qquad\qquad x^{(i)}_{j,j^{\prime}}\in[0,1]\qquad\text{for all $j,j^{\prime}$}$
@@
As described in [1], this relaxation can be solved in polynomial time. Furthermore, given an integral solution $X^{(i)}_{j}\in\{0,1\}^{\mathcal{J}}$, there is a corresponding SDP solution defined by $x^{(i)}_{j,j^{\prime}}=X^{(i)}_{j}X^{(i)}_{j^{\prime}}$; note then that $\mathbf{X}^{(i)}=(1,x^{(i)}_{1},\dots,x^{(i)}_{n})\times(1,x^{(i)}_{1},\dots,x^{(i)}_{n})^{\top}$, which is an outer product and hence positive-semidefinite. Our goal is to convert the fractional assignments $x^{(i)}_{j}$ into integral assignments $X^{(i)}_{j}$. We quote the following key results of [1] concerning the semidefinite program.
@@
We define $\mathcal{J}^{*}$ to be the set of jobs $j$ with $\sigma_{j}\geq\sigma_{j^{*}}$. As a point of notation, any sum of the form $\sum_{j}$ should be taken to range over $j\in\mathcal{J}^{*}$ unless stated otherwise. Likewise, in a sum of the form $\sum_{j,j^{\prime}}$, we view $j,j^{\prime}$ as an ordered pair of jobs in $\mathcal{J}^{*}$; there are separate summands for $j,j^{\prime}$ and for $j^{\prime},j$ and we also allow $j=j^{\prime}$.
For each class $\mathcal{P}_{k}$ we define $\mathcal{P}^{*}_{k}=\mathcal{P}_{k}\cap\mathcal{J}^{*}$ and for each cluster $\mathcal{C}_{k,\ell}$, we define $\mathcal{C}^{*}_{k,\ell}=\mathcal{C}_{k,\ell}\cap\mathcal{J}^{*}$. Within each class $\mathcal{P}_{k}$, the final opened cluster $\mathcal{C}^{*}_{k,\ell}$ after processing job $j^{*}$ is called the leftover cluster for $k$, and denoted by $\mathcal{C}^{*}_{k,\text{left}}=\mathcal{C}^{*}_{k,\ell}$.
@@
For any vector $y\in[0,1]^{\mathcal{J}^{*}}$ there holds
@@
We remark that Theorem 20 was shown in [1] for an integral vector $y\in\{0,1\}^{\mathcal{J}^{*}}$, but we will need the slightly generalized result for our analysis.
@@ 5 Determining the approximation ratio
We now begin the computation for the approximation ratio. The analysis has three main steps. First, we compute an upper bound on the expected value of the solution returned by the rounding algorithm. Second, we compute a lower bound on the objective function of the relaxation. Finally, we combine these two estimates.
@@ 5.1 The algorithm upper-bound
where recall that
$\phi_{j,j^{\prime}}=\frac{((1-\rho_{j})^{1-1/x_{j}}-1)((1-\rho_{j^{\prime}})^{1-1/x_{j^{\prime}}}-1)}{(1-\rho_{j})^{1-1/x_{j}}(1-\rho_{j^{\prime}})^{1-1/x_{j^{\prime}}}+\rho_{j}+\rho_{j^{\prime}}-1}$
@@
The main focus of the analysis is to show that $B_{k,\ell}$ is large on average. For intuition, keep in mind the extremal case when all the untruncated jobs have infinitesimal mass (i.e. $x_{j}^{2}\approx 0$). In this case, the jobs in each cluster $\mathcal{C}^{*}_{k,\ell}$ have $\phi_{j,j^{\prime}}=\frac{\mathrm{e}^{1/\lambda}-1}{\mathrm{e}^{1/\lambda}+1}$ where $\lambda=\sum_{j\in\mathcal{C}^{*}_{k,\ell}}\tilde{\rho}_{i,j}$. The truncated job (if any) requires some special analysis. We also need to show that non-infinitesimal mass can only increase the value $B_{k,\ell}$.
@@
For a cluster $k,\ell$, let $\mathcal{U}$ be the set of untruncated jobs in $\mathcal{C}^{*}_{k,\ell}$ and define parameters $\lambda,a,r,s$ as follows:
$\lambda=\sum_{j\in\mathcal{C}^{*}_{k,\ell}}\tilde{\rho}_{j},\quad a=\frac{\mathrm{e}^{1/\lambda}-1}{\mathrm{e}^{1/\lambda}+1},\quad r=\sum_{j\in\mathcal{U}}x_{j},\quad s=\sum_{j\in\mathcal{U}}x_{j}H_{j}$
There holds:
$\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{U}:j\neq j^{\prime}\end{subarray}}\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}H_{j}H_{j^{\prime}}\geq\sum_{j\in\mathcal{U}}x_{j}H_{j}(a-c_{0}x_{j})(s-x_{j})\quad\text{for constant $c_{0}:=0.17556$}$
@@
Each untruncated job $j$ has $\rho_{j}=\tilde{\rho}_{j}/\lambda=x_{j}/\lambda$ where $\lambda\leq\tau\leq 3/4$. By Lemma 14, any pair $j,j^{\prime}\in\mathcal{U}$ has $\phi_{j,j^{\prime}}\geq a-b(x_{j}+x_{j^{\prime}})$ for $b=0.57\max\{0,\lambda-0.45\}\leq 0.57(\tau-0.45)=c_{0}/2$. So:
@@
[…]
@@
Finally, since $a\geq\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}\geq 0.68$ and $H_{j}\geq 1$, we have $(a-c_{0}x_{j})(s-x_{j}H_{j})\geq(a-c_{0}x_{j})(s-x_{j})$. ∎
Lemma 25.
For a cluster $k,\ell$, define parameters $\lambda,a,r,s$ as in Lemma 24.
If $\mathcal{C}^{*}_{k,\ell}$ has no truncated jobs, then $B_{k,\ell}\geq s/2\cdot\min_{x\in[0,r]}\bigl{(}x+(a-c_{0}x)(s-x)\bigr{)}$.
If $\mathcal{C}^{*}_{k,\ell}$ has a truncated job $j_{\text{tr}}$, then define $y=x_{j_{\text{tr}}},d=x_{j_{\text{tr}}}H_{j_{\text{tr}}}$ and we have:
$B_{k,\ell}\geq d^{2}/2+s/2\cdot\min_{x\in[0,r]}\Bigl{(}x+(a-c_{0}x)(s-x)+\frac{2d((1-x/\tau)^{1-1/x}-1)((r/\tau)^{1-1/y}-1)}{(1-x/\tau)^{1-1/x}(r/\tau)^{1-1/y}-(r-x)/\tau}\Bigr{)}$
Proof.
Let $\mathcal{T}$ denote the set of truncated jobs in $\mathcal{C}^{*}_{k,\ell}$; note that $|\mathcal{T}|\leq 1$. By Lemma 24, can decompose $B_{k,\ell}$ as follows:
$2B_{k,\ell}$ $=\sum_{j_{\text{tr}}\in\mathcal{T}}x_{j_{\text{tr}}}^{2}H_{j_{\text{tr}}}^{2}+\sum_{j\in\mathcal{U}}x_{j}^{2}H_{j}^{2}+\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{U}:j\neq j^{\prime}\end{subarray}}\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}H_{j}H_{j^{\prime}}+2\sum_{\begin{subarray}{c}j\in\mathcal{U},j_{\text{tr}}\in\mathcal{T}\end{subarray}}\phi_{j,j_{\text{tr}}}x_{j}x_{j_{\text{tr}}}H_{j}H_{j_{\text{tr}}}$
$\geq\sum_{j_{\text{tr}}\in\mathcal{T}}x_{j_{\text{tr}}}^{2}H_{j_{\text{tr}}}^{2}+\sum_{j\in\mathcal{U}}x_{j}H_{j}\Bigl{(}x_{j}+(a-c_{0}x_{j})(s-x_{j})+2\sum_{j_{\text{tr}}\in\mathcal{T}}\phi_{j,j_{\text{tr}}}x_{j_{\text{tr}}}H_{j_{\text{tr}}}\Bigr{)}$
If $\mathcal{T}=\emptyset$, then $2B_{k,\ell}\geq\sum_{j\in\mathcal{U}}x_{j}H_{j}(x_{j}+(a-c_{0}x_{j})(s-x_{j}))$. Furthermore, $x_{j}\in[0,r]$, so we lower-bound this in turn by:
$2B_{k,\ell}\geq\sum_{j\in\mathcal{U}}x_{j}H_{j}\min_{x\in[0,r]}\bigl{(}x+(a-c_{0}x)(s-x)\bigr{)}=s\min_{x\in[0,r]}\bigl{(}x+(a-c_{0}x)(s-x)\bigr{)}$
as desired. Next, if $\mathcal{C}^{*}_{k,\ell}$ has a truncated job $j_{\text{tr}}$ with $\rho_{j_{\text{tr}}}=1-r/\tau$, then for any job $j\in\mathcal{U}$ we have
$\phi_{j,j_{\text{tr}}}=\frac{((1-x_{j}/\tau)^{1-1/x_{j}}-1)((r/\tau)^{1-1/y}-1)}{(1-x_{j}/\tau)^{1-1/x_{j}}(r/\tau)^{1-1/y}-(r-x_{j})/\tau}.$
and we get:
$2B_{k,\ell}\geq d^{2}+\sum_{j\in\mathcal{U}}x_{j}H_{j}\Bigl{(}x_{j}+(a-c_{0}x_{j})(s-x_{j})+\frac{2d((1-x_{j}/\tau)^{1-1/x_{j}}-1)((r/\tau)^{1-1/y}-1)}{(1-x_{j}/\tau)^{1-1/x_{j}}(r/\tau)^{1-1/y}-(r-x_{j})/\tau}\Bigr{)}$
and the result follows again since $x_{j}\in[0,r]$. ∎
@@
Let $\lambda,a,r,s$ be as in Lemma 24, and define $V=\sum_{j\in\mathcal{C}^{*}_{k,\ell}}x_{j}(H_{j}-\kappa)$.
If $\mathcal{C}^{*}_{k,\ell}$ has no truncated job, then $\lambda=r\in[\theta,\tau]$ and $V=s-\kappa r$. So $a=\frac{\mathrm{e}^{1/r}-1}{\mathrm{e}^{1/r}+1}$ and Lemma 25 gives
$\frac{B_{k,\ell}}{V}\geq\min_{\begin{subarray}{c}r\in[\theta,\tau],x\in[0,r]\\ s\in[r,\pi r]\end{subarray}}\frac{s\bigl{(}x+(\frac{\mathrm{e}^{1/r}-1}{\mathrm{e}^{1/r}+1}-c_{0}x)(s-x)\bigr{)}}{2(s-\kappa r)}$
@@
If $\mathcal{C}^{*}_{k,\ell}$ has a truncated job $j_{\text{tr}}$, then let $y=x_{j_{\text{tr}}}\geq\tau-r$ and $d=x_{j_{\text{tr}}}H_{j_{\text{tr}}}$. So $V=s+d-\kappa(r+y)$ and $\lambda=\tau$ and $a=\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}$. Lemma 25 gives:
$\frac{B_{k,\ell}}{V}\geq\min_{\begin{subarray}{c}r\in[0,\theta],x\in[0,r]\\ y\in[\tau-r,1]\\ s\in[r,\pi r],d\in[y,\pi y]\end{subarray}}\frac{d^{2}+s\bigl{(}x+(\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}-c_{0}x)(s-x)+2d\frac{((1-x/\tau)^{1-1/x}-1)((r/\tau)^{1-1/y}-1)}{(1-x/\tau)^{1-1/x}(r/\tau)^{1-1/y}-(r-x)/\tau}\bigr{)}}{2(s+d-\kappa(r+y))}$
@@
Then $R_{k}\in[0,\theta],T_{k}\in[1,\pi]$ and $B_{k,\ell}\geq R_{k}\cdot\min\{c_{1}(T_{k}-\kappa),c_{2}R_{k}T_{k}^{2}\}$ for constant $c_{2}:=0.33965$.
(If $R_{k}=0$, we may set $T_{k}$ to an arbitrary value in $[1,\pi]$ and the result still holds.)
@@
The cluster $\mathcal{C}^{*}_{k,\text{left}}$ has no truncated jobs and has $R_{k}\in[0,\theta]$, as otherwise it would be closed after adding its final job. In the language of Lemma 24, we have $r=R_{k},s=S_{k}=R_{k}T_{k}$, and $a\geq\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}\approx 0.67930078\geq 2c_{2}=0.6793$.
Lemma 25 gives $B_{k,\ell}\geq s/2\cdot\min_{x\in[0,r]}(x+(2c_{2}-c_{0}x)(s-x)\bigr{)}$. Thus, removing a common factor of $r$ and letting $t=T_{k}$, it suffices to show that
@@
[…]
@@ 5.2 The algorithm lower-bound
$A^{2}\leq\sum_{k}P_{k}^{2}\sum_{j\in\mathcal{C}_{k}}x_{j}g_{k}(H_{j})\Bigl{(}c_{4}+R_{k}\cdot\frac{T_{k}+\delta}{H_{j}+\delta}\cdot g_{k}(H_{j})\Bigr{)}$ for constant $c_{4}:=0.035898$.
@@
${\mathbb{E}}[A^{2}]\leq\gamma{\mathbb{E}}[D]+c_{5}Q\qquad\text{for constant $c_{5}:=0.004562$}$
@@
$A^{2}-\gamma D$ $\leq\sum_{k}P_{k}^{2}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}\Bigl{(}g_{k}(H_{j})\bigl{(}c_{4}+\frac{R_{k}(T_{k}+\delta)g_{k}(H_{j})}{H_{j}+\delta}\bigr{)}-\gamma f(R_{k},T_{k})\frac{H_{j}+\delta}{T_{k}+\delta}\Bigr{)}$ (6)
@@
$F(h)=\max_{\begin{subarray}{c}r\in[0,\theta],t\in[1,\pi]\end{subarray}}\ \ \ g(r,t,h)\bigl{(}c_{4}+\frac{rg(r,t,h)(t+\delta)}{h+\delta}\bigr{)}-\frac{\gamma f(r,t)(h+\delta)}{t+\delta}$
@@
[…]

### v8

@@ 4.1 The SDP relaxation
Here, we provide a brief summary of the SDP relaxation and its properties. See [2] for more details.
@@
[…]
@@
As described in [2], this relaxation can be solved in polynomial time. Furthermore, given an integral solution $X^{(i)}_{j}\in\{0,1\}^{\mathcal{J}}$, there is a corresponding SDP solution defined by $x^{(i)}_{j,j^{\prime}}=X^{(i)}_{j}X^{(i)}_{j^{\prime}}$. Our goal is to convert the fractional assignments $x^{(i)}_{j}$ into integral assignments $X^{(i)}_{j}$. We quote the following key results of [2] concerning the semidefinite program.
@@
We define $\mathcal{J}^{*}$ to be the set of jobs $j$ with $x_{j}>0$ and $\sigma_{j}\geq\sigma_{j^{*}}$. As a point of notation, any sum of the form $\sum_{j}$ should be taken to range over $j\in\mathcal{J}^{*}$ unless stated otherwise. Likewise, in a sum of the form $\sum_{j,j^{\prime}}$, we view $j,j^{\prime}$ as an ordered pair of jobs in $\mathcal{J}^{*}$; there are separate summands for $j,j^{\prime}$ and for $j^{\prime},j$ and we also allow $j=j^{\prime}$.
For each class $\mathcal{P}_{k}$ we define $\mathcal{P}^{*}_{k}=\mathcal{P}_{k}\cap\mathcal{J}^{*}$ and for each cluster $\mathcal{C}_{k,\ell}$, we define $\mathcal{C}^{*}_{k,\ell}=\mathcal{C}_{k,\ell}\cap\mathcal{J}^{*}$. Within each class $\mathcal{P}_{k}$, the final opened cluster $\mathcal{C}^{*}_{k,\ell}$ after processing job $j^{*}$ is called the leftover cluster for $k$, and denoted by $\mathcal{C}^{*}_{k,\text{left}}$.
@@
For any vector $y\in[0,1]^{\mathcal{J}^{*}}$ there holds [^4] Theorem 20 was shown in [2] for an integral vector $y\in\{0,1\}^{\mathcal{J}^{*}}$, but we will need the slightly generalized result for our analysis.
@@ 5 Determining the approximation ratio
We now begin the arduous computation of the approximation ratio. The analysis has three main steps. First, we compute an upper bound on the expected value of the solution returned by the rounding algorithm. Second, we compute a lower bound on the objective function of the relaxation. Finally, we combine these two estimates.
@@ 5.1 The algorithm upper-bound
where recall that $\phi_{j,j^{\prime}}=\Phi(x_{j},x_{j^{\prime}};\rho_{j},\rho_{j^{\prime}}).$
@@
The main focus of the analysis is to show that $B_{k,\ell}$ is large on average. For intuition, keep in mind the extremal case when all untruncated jobs have infinitesimal mass (i.e. $x_{j}^{2}\approx 0$). The truncated jobs require some special analysis. We also need to show that non-infinitesimal mass can only increase the value $B_{k,\ell}$.
@@
For a cluster $k,\ell$, let $\mathcal{U},\mathcal{T}$ be the set of untruncated jobs and truncated jobs respectively in $\mathcal{C}^{*}_{k,\ell}$ and define parameters as follows:
$\lambda=\sum_{j\in\mathcal{C}^{*}_{k,\ell}}\tilde{\rho}_{j},\quad a=\frac{\mathrm{e}^{1/\lambda}-1}{\mathrm{e}^{1/\lambda}+1},\quad r=\sum_{j\in\mathcal{U}}x_{j},\quad s=\sum_{j\in\mathcal{U}}x_{j}H_{j},\quad y=\sum_{j\in\mathcal{T}}x_{j},\quad d=\sum_{j\in\mathcal{T}}x_{j}H_{j}$
We have:
$B_{k,\ell}\geq d^{2}/2+s/2\cdot\inf_{x\in(0,r]}\Bigl(x+(a-c_{0}x)(s-x)+2d\Phi(x,y;x/\tau,1-r/\tau)\Bigr)$
for constant $c_{0}:=0.17556$.
@@
We first claim that there holds:
$\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{U}:j\neq j^{\prime}\end{subarray}}\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}H_{j}H_{j^{\prime}}\geq\sum_{j\in\mathcal{U}}x_{j}H_{j}(a-c_{0}x_{j})(s-x_{j})\quad$ (5)
For, each untruncated job $j$ has $\rho_{j}=\tilde{\rho}_{j}/\lambda=x_{j}/\lambda$ where $\lambda\leq\tau\leq 3/4$. Because we only add jobs with non-zero mass to clusters, it also must have $x_{j}>0$. By Lemma 14, any pair $j,j^{\prime}\in\mathcal{U}$ thus has $\phi_{j,j^{\prime}}\geq a-b(x_{j}+x_{j^{\prime}})$ for $b=0.57\max\{0,\lambda-0.45\}\leq 0.57(\tau-0.45)=c_{0}/2$. So:
@@
$\qquad=\sum_{j,j^{\prime}\in\mathcal{U}:j\neq j^{\prime}}\bigl(a-c_{0}x_{j}\bigr)x_{j}x_{j^{\prime}}H_{j}H_{j^{\prime}}\qquad\text{(by symmetry)}$
@@
Finally, since $a\geq\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}\geq 0.68$ and $H_{j}\geq 1$, we have $(a-c_{0}x_{j})(s-x_{j}H_{j})\geq(a-c_{0}x_{j})(s-x_{j})$, which establishes the bound (5).
Now, noting that $|\mathcal{T}|\leq 1$, we can decompose $B_{k,\ell}$ as follows:
$2B_{k,\ell}$ $=\sum_{t\in\mathcal{T}}x_{t}^{2}H_{t}^{2}+\sum_{j\in\mathcal{U}}x_{j}^{2}H_{j}^{2}+\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{U}:j\neq j^{\prime}\end{subarray}}\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}H_{j}H_{j^{\prime}}+2\sum_{\begin{subarray}{c}j\in\mathcal{U},t\in\mathcal{T}\end{subarray}}\phi_{j,t}x_{j}x_{t}H_{j}H_{t}$
$\geq\sum_{t\in\mathcal{T}}x_{t}^{2}H_{t}^{2}+\sum_{j\in\mathcal{U}}x_{j}H_{j}\Bigl(x_{j}+(a-c_{0}x_{j})(s-x_{j})+2\sum_{t\in\mathcal{T}}\phi_{j,t}x_{t}H_{t}\Bigr)$
If $\mathcal{C}^{*}_{k,\ell}$ has a truncated job $t$ with $\rho_{t}=1-r/\tau$, then for any job $j\in\mathcal{U}$ we have $\phi_{j,t}=\Phi(x_{j},y;x_{j}/\tau,1-r/\tau)$ and we get:
$2B_{k,\ell}\geq d^{2}+\sum_{j\in\mathcal{U}}x_{j}H_{j}\bigl(x_{j}+(a-c_{0}x_{j})(s-x_{j})+2d\Phi(x_{j},y;x_{j}/\tau,1-r/\tau)\bigr)$
and the result follows since $x_{j}\in(0,r]$. ∎
@@
Let $\lambda,a,r,s,y,d$ be as in Lemma 24, and define $V=\sum_{j\in\mathcal{C}^{*}_{k,\ell}}x_{j}(H_{j}-\kappa)$.
If $\mathcal{C}^{*}_{k,\ell}$ has no truncated job, then $y=d=0$ and $\lambda=r\in[\theta,\tau]$ and $V=s-\kappa r$. So $a=\frac{\mathrm{e}^{1/r}-1}{\mathrm{e}^{1/r}+1}$ and Lemma 24 gives
$\frac{B_{k,\ell}}{V}\geq\inf_{\begin{subarray}{c}r\in[\theta,\tau],x\in(0,r]\\ s\in[r,\pi r]\end{subarray}}\frac{s\bigl(x+(\frac{\mathrm{e}^{1/r}-1}{\mathrm{e}^{1/r}+1}-c_{0}x)(s-x)\bigr)}{2(s-\kappa r)}$
@@
If $\mathcal{C}^{*}_{k,\ell}$ has a truncated job, then $y\geq\tau-r$, and so $V=s+d-\kappa(r+y)$ and $\lambda=\tau$ and $a=\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}$. Lemma 24 gives:
$\frac{B_{k,\ell}}{V}\geq\inf_{\begin{subarray}{c}r\in[0,\theta],x\in(0,r]\\ y\in[\tau-r,1]\\ s\in[r,\pi r],d\in[y,\pi y]\end{subarray}}\frac{d^{2}+s\bigl(x+(\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}-c_{0}x)(s-x)+2d\Phi(x,y;x/\tau,1-r/\tau)\bigr)}{2(s+d-\kappa(r+y))}$
@@
Then $R_{k}\in[0,\theta],T_{k}\in[1,\pi]$ and
$B_{k,\ell}\geq R_{k}\cdot\min\{c_{1}(T_{k}-\kappa),c_{2}R_{k}T_{k}^{2}\}\qquad\text{for constant $c_{2}:=0.33965$}$
(If $R_{k}=0$, we may set $T_{k}$ to an arbitrary value in $1,\pi]$ and the result still holds.)
@@
The cluster $\mathcal{C}^{*}_{k,\text{left}}$ has no truncated jobs and has $R_{k}\in[0,\theta]$, as otherwise it would be closed after adding its final job. In the language of Lemma 24, we have $r=R_{k},s=S_{k}=R_{k}T_{k}$, and $y=d=0$, and $a\geq\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}\approx 0.67930078\geq 2c_{2}=0.6793$.
Lemma 24 gives $B_{k,\ell}\geq s/2\cdot\inf_{x\in(0,r]}(x+(2c_{2}-c_{0}x)(s-x)\bigr)$. Thus, removing a common factor of $r$ and letting $t=T_{k}$, it suffices to show that
@@
${\mathbb{E}}[(H_{j}-\kappa)/H_{j}^{2}]=\frac{1}{\log\pi}\int_{h=1}^{\pi}\frac{h-\kappa}{h^{3}}\ \mathrm{d}h=\frac{\kappa-2\pi+2\pi^{2}-\kappa\pi^{2}}{2\pi^{2}\log\pi}$
@@ 5.2 The algorithm lower-bound
$A^{2}\leq\sum_{k}P_{k}^{2}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}g_{k}(H_{j})\Bigl(c_{4}+R_{k}\cdot\frac{T_{k}+\delta}{H_{j}+\delta}\cdot g_{k}(H_{j})\Bigr)$ for constant $c_{4}:=0.035898$.
@@
${\mathbb{E}}[A^{2}]\leq\gamma{\mathbb{E}}[D]+c_{5}Q\qquad\text{for constant $c_{5}:=0.004562$}.$
@@
$A^{2}-\gamma D$ $\leq\sum_{k}P_{k}^{2}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}\Bigl(g_{k}(H_{j})\bigl(c_{4}+\frac{R_{k}(T_{k}+\delta)g_{k}(H_{j})}{H_{j}+\delta}\bigr)-\gamma\frac{H_{j}+\delta}{T_{k}+\delta}f(R_{k},T_{k})\Bigr)$ (7)
@@
$F(h)=\max_{\begin{subarray}{c}r\in[0,\theta],t\in[1,\pi]\end{subarray}}\ \ \ g(r,t,h)\bigl(c_{4}+\frac{r(t+\delta)g(r,t,h)}{h+\delta}\bigr)-\frac{\gamma f(r,t)(h+\delta)}{t+\delta}$
@@
By Observation 17, each item $j$ has ${\mathbb{E}}[F(H_{j})/H_{j}^{2}]=\frac{1}{\log\pi}\int_{h=1}^{\pi}F(h)/h^{3}\ \mathrm{d}h$, giving us
${\mathbb{E}}[A^{2}]$ $\leq\gamma{\mathbb{E}}[D]+\Bigl(\frac{1}{\log\pi}\int_{h=1}^{\pi}\frac{F(h)}{h^{3}}\ \mathrm{d}h\Bigr)Q$
We show in Appendix B that $c_{5}\geq\frac{1}{\log\pi}\int_{h=1}^{\pi}\frac{F(h)}{h^{3}}\ \mathrm{d}h$, which yields the claimed result. ∎